Dilly Farm Supply is located in a small town in the rural west. Data regarding the store's operations follow:

  • Sales are budgeted at $251,000 for November, $291,000 for December, and $218,000 for January.
  • Collections are expected to be 60% in the month of sale and 40% in the month following the sale.
  • The cost of goods sold is 75% of sales.
  • The company desires to have an ending merchandise inventory at the end of each month equal to 90% of the next month's cost of goods sold. Payment for merchandise is made in the month following the purchase.
  • Other monthly expenses to be paid in cash are $20,000.
  • Monthly depreciation is $25,000.
  • Ignore taxes.
Balance Sheet
October 31
Assets
Cash $ 26,000
Accounts receivable 81,000
Merchandise inventory 169,425
Property, plant and equipment, net of $624,000 accumulated depreciation 946,000
Total assets $ 1,222,425
Liabilities and Stockholders' Equity
Accounts payable $ 247,000
Common stock 748,000
Retained earnings 227,425
Total liabilities and stockholders' equity $ 1,222,425

The difference between cash receipts and cash disbursements for December would be:

Brewer 8e Rechecks 2019-26-08

Multiple Choice

  • $20,750

  • $13,750

  • $39,750
Answer #1
Cost of goods sold for November 188250 =251000*75%
Add: Desired ending inventory 196425 =291000*75%*90%
Less: Beginning inventory -169425 =251000*75%*90%
Purchases for November 215250
Collection from November sales 100400 =251000*40%
Collection from December sales 174600 =291000*60%
Cash receipts for December 275000
Payment for November purchases 215250
Other monthly expenses 20000
Cash disbursements for December 235250
Cash receipts for December 275000
Less: Cash disbursements for December 235250
Difference between cash receipts
and cash disbursements
39750
Option C $39,750 is correct